# Improving Diffusion-based Data Augmentation with Inversion Spherical Interpolation

### v1

@@ Abstract
Data Augmentation (DA), i.e., synthesizing faithful and diverse samples to expand the original training set, is a prevalent and effective strategy to improve various visual recognition tasks. With the powerful image generation ability, diffusion-based DA has shown strong performance gains on different benchmarks. In this paper, we analyze today’s diffusion-based DA methods, and argue that they cannot take account of both faithfulness and diversity, which are two critical keys for generating high-quality samples and boosting final classification performance. To this end, we propose a novel Diffusion-based Inversion Interpolation DA method: Diff-II. Specifically, Diff-II consists of three main steps: 1) Category concepts learning: Learning concept embeddings for each category. 2) Inversion interpolation: Calculating the inversion for each image, and conducting spherical interpolation for two randomly sampled inversions from the same category. 3) Two-stage denoising: Using different prompts to generate synthesized images in a coarse-to-fine manner. Extensive experiments on multiple image classification tasks (e.g., few-shot, long-tailed, and out-of-distribution classification) have demonstrated its effectiveness over state-of-the-art diffusion-based DA methods.
Figure 1: Based on given training images, data augmentation (DA) aims to generate new faithful and diverse synthetic images. (a) These synthetic images are faithful but not diverse. (b) These synthetic images are diverse but not faithful. (c) These synthetic images are both faithful and diverse.
@@ 1 Introduction
Today’s visual recognition models can even outperform us humans with sufficient training samples. However, in many different real-world scenarios, it is not easy to collect adequate training data for some categories. For example, since the occurrence frequency of various categories in nature follows a long-tailed distribution, there are many rare categories with only limited samples (O’Hagan & Forster, 2004; Van Horn & Perona, 2017). To mitigate this data scarcity issue, a prevalent and effective solution is Data Augmentation (DA). Based on an original training set with limited samples, DA aims to generate more synthetic samples to expand the training set.
@@
For DA methods, there are two critical indexes: faithfulness and diversity (Sajjadi et al., 2018). They can not only evaluate the quality of synthesized samples, but also influence the final recognition performance. As shown in Figure 1, faithfulness indicates that the synthetic samples need to retain the characteristics of the corresponding category (c.f., Figure 1 (a)), i.e., the faithfulness confirms that the model learns from correct category knowledge. Diversity indicates that the synthetic samples should have different contexts from the original training set and each other (c.f., Figure 1 (b)), i.e., the diversity ensures that the model learns the invariable characteristics of the category by seeing diverse samples.
@@
Figure 2: a) Intra-category DA: Given a reference image (from the original set), it adds some noise and denoises with a prompt containing the same category concept (e.g., concept “[A]” for category A image). (b) Inter-category DA: Different from Intra-category DA, it denoises with a prompt containing a different category concept (e.g., concept “[B]” for category A image). (c) Ours: It first calculates the inversion for each image, and conducts spherical interpolation for two images of the same category. Then, it denoises in a two-stage manner with different prompts.
@@
With the photo-realistic image generation ability of today’s diffusion models (Ho et al., 2020; Nichol & Dhariwal, 2021), a surge of diffusion-based DA methods has dominated the image classification task. Typically, diffusion-based DA methods reformulate image augmentation as an image editing task, which consists of two steps: 1) Noising Step: They first randomly sample an image from the original training set as a reference image and then add some noise to the reference image. 2) Denoising Step: They then gradually denoise this noisy reference image conditioned on a category-specific prompt. After the two steps, a new synthesized training image was generated. Following this framework, the pioneer diffusion-based DA work (He et al., 2022) directly uses a hand-crafted template containing the reference image’s category label as the prompt (i.e., intra-category denoising). These handcrafted prompts work well on general datasets with a broad spectrum of category concepts (e.g., CIFAR-10 (Krizhevsky et al., 2009)). However, these few words (with only category name) can not guide the diffusion models to generate images with specific and detailed characteristics, especially for datasets with fine-grained categories (e.g., Stanford Cars (Krause et al., 2013)).
@@
To further enhance the generalization ability, subsequent diffusion-based DA methods try to improve the quality of synthesized samples from the two key characteristics. Specifically, to improve faithfulness, Trabucco et al. (2023) replace category labels with more fine-grained learned category concepts. As shown in Figure 2(a), they first learn a specific embedding “[A]” for “category A” bird, and then replace the fixed category name with the learned concept in the prompt. These learnable prompts can somewhat preserve fine-grained details for different categories. However, the fixed combination of a reference image and its corresponding category concept always results in similar synthetic samples (i.e., limited diversity). On the other side, to improve diversity, Wang et al. (2024) use prompts containing different category concepts (e.g., “[B]”) from the reference image (i.e., inter-category denoising). This operation can generate images with “intermediate” semantics between two different categories. However, it inherently introduces another challenging problem to obtain an “accurate” soft label for each synthetic image, which affects faithfulness to some extent (c.f., Figure 2(b)). Based on these above discussions, we can observe that: current state-of-the-art diffusion-based DA methods cannot take account of both faithfulness and diversity, which results in limited improvements on the generalization ability of downstream classifiers.
@@
In this paper, we propose a simple yet effective Diffusion-based Inversion Interpolation method: Diff-II, which can generate both faithful and diverse augmented images. As shown in Figure 2(c), Diff-II consists of three steps: 1) Category Concepts Learning: To generate faithful images, we learn a specific embedding for each category (e.g., “[A]” for category A) by reconstructing the images of the original training set. 2) Inversion Interpolation: To improve diversity while maintaining faithfulness, we calculate the inversion [^1] In the image generation field, the inversion refers to a latent representation that can be used to reconstruct the corresponding original image by the generative model. for each image of the original training set. Then, we sample two inversions from the same category and conduct interpolation. The interpolation result corresponds to a subsequent high-quality synthetic image. 3) Two-stage Denoising: To further improve the diversity, we prepare some suffixes [^2] More details are left in the appendix. (e.g., “flying over water”, “standing on a tree branch”) that can summarize the high-frequency context patterns of the original training set. Then, we split the denoising process into two stages by timesteps. In the first stage, we denoise the interpolation results guided by a prompt containing the learned category concept and a randomly sampled suffix, e.g., “a photo of a [A] bird [suffix].” This design can inject perturbation into the early-timestep generation of context and finally contributes to diversity. In the second stage, we replace the prompt with “a photo of a [A] bird” to refine the character details of the category concept.
@@
To be specific, we first utilize some parameter-efficient fine-tuning methods (e.g., low-rank adaptation (Hu et al., 2021) and textual inversion (Gal et al., 2022)) to learn the concept embedding for each category. Then, we acquire the DDIM inversion (Song et al., 2020) for each image from the original set conditioned on the learned concept. After that, we randomly sample two inversions within one category as one pair and conduct interpolation (both inward and outward) with random strengths. To align the distribution of interpolation results with standard normal distribution, we conduct spherical interpolation (Shoemake, 1985). Since each pair of images used for inversion interpolation belongs to the same concept, their interpolations will highly maintain the semantic consistency of this concept (i.e., it ensures faithfulness). Meanwhile, since both images have different contexts, the interpolations will produce an image with a new context (i.e., it ensures diversity). Finally, we set a split ratio to divide the whole denoising timesteps into two stages. In the first stage, we use a prompt containing the learned concept and a randomly sampled suffix2 to generate noisy images with diverse contexts (e.g., layout and gesture). In the second stage, we remove the suffix to refine the character details of the category concept. By adjusting the split ratio, we can control the trade-off between faithfulness and diversity. To extract all suffixes, we first utilize a pretrained vision-language model to extract all captions of the original training set, and then leverage a large language model to summarize them into a few suffixes.
@@
  • We use a unified view to analyze existing diffusion-based DA methods, we argue that they can not take account of both faithfulness and diversity well, which leads to limited improvements.
  • We propose Diff-II, a simple yet effective diffusion-based DA method, that leverages the inversion interpolation and two-stage denoising to generate faithful and diverse images.
  • We conduct comprehensive experiments on three tasks. Our state-of-the-art performance verifies that Diff-II can achieve effective data augmentation by generating high-quality samples.
@@ 2 Related Work
Diffusion-Based DA. Currently, there are two main paradigms of Diffusion-based DA: 1) Latent Perturbation (Zhou et al., 2023; Fu et al., 2024; Zhang et al., 2024) generate samples by perturbating latent codes in the latent space. Although these methods can generate diverse samples, due to the uncontrollable perturbation direction, the generated results sometimes deviate from the domain of the original dataset. Therefore, they heavily rely on extra over-sampling and filtering steps. 2) Image Editing (He et al., 2022; Trabucco et al., 2023; Dunlap et al., 2023; Wang et al., 2024) reformulate data augmentation as an image editing task (Meng et al., 2021; Lu et al., 2023). However, due to the limitations of the editing paradigm, it is difficult for them to take into account both the faithfulness and diversity of the synthetic samples. To tackle this problem, our work proposes to generate new images by interpolating the inversions.
Interpolation-Based Data Augmentation. For time series and text data, interpolation is a common approach for DA. Chen et al. (2022) incorporate a two-stage interpolation in the hidden space to improve the text classification models. Oh et al. (2020) propose to augment time-series data by interpolation on original data. In the computer vision community, there are some studies (DeVries & Taylor, 2017; Zhou et al., 2023) work on interpolation-based DA for image classification. However, how to combine the excellent generation ability of diffusion models and interpolation operation to obtain high-quality synthetic samples remains an important challenge. To solve this problem, we propose Diff-II, an efficient diffusion-based DA framework with inversion interpolation.
@@ 3 Method
Figure 3: Pipeline of Diff-II. (1) Concept Learning: Learning accurate concepts for each category. (2) Inversion Interpolation: Calculating DDIM inversion for each image conditioned on the learned concept. Then, randomly sampling a pair and conducting random spherical interpolation. (3) Two-stage Denoising: Denosing the interpolation results in a two-stage manner with different prompts.
@@
1) Category Concepts Learning (Sec. 3.1): We first set $n$ learnable token embeddings for each category, and insert some learnable low-rank matrixes into the pretrained diffusion U-Net. By reconstructing the noised image of the original training set $\mathcal{O}$, we learn the accurate concept for each category. We denote the tokens of the $i_{th}$ category concept as $\{[V_{j}^{i}]\}_{j=1}^{n}$.
2) Inversion Interpolation (Sec. 3.2): Take the $i_{th}$ category as an example, we form a prompt: “a photo of a $[V_{1}^{i}]$ $[V_{2}^{i}]$$[V_{n}^{i}]$ [metaclass]2”. The “[metaclass]” is the theme of the corresponding dataset, e.g.“bird” is the “[metaclass]” for dataset CUB (Wah et al., 2011). Then, we calculate the DDIM inversion $I_{j}^{i}$ for each training sample $X_{j}^{i}\in\mathcal{O}^{i}$ conditioned on this prompt. All these inversions (from $\mathcal{O}^{i}$) made up the inversion pool $\mathcal{I}^{i}=\{I_{j}^{i}\}_{j=1}^{N_{i}}$ (c.f., Sec. 3.2.1). After that, we randomly sample two inversions ($I_{a}^{i}$, $I_{b}^{i}$) from $\mathcal{I}^{i}$ and conduct random spherical interpolation on this pair (c.f., Sec. 3.2.2). The interpolation result is denoted as $Z$. We repeat the sampling and interpolation then collect all interpolation results into $\mathcal{Z}^{i}$.
3) Two-stage Denoising (Sec. 3.3): Given an interpolation $Z\in\mathcal{Z}^{i}$, we denoise it as the initial noise in two stages. The main difference between the two stages is the prompt used. In the first stage, we use a suffixed prompt: “a photo of a $[V_{1}^{i}]$ $[V_{2}^{i}]$$[V_{n}^{i}]$ [metaclass] [suffix]”. In the second stage, we use a plain prompt: “a photo of a $[V_{1}^{i}]$ $[V_{2}^{i}]$$[V_{n}^{i}]$ [metaclass]”. Repeat two-stage denoising for each $Z\in\mathcal{Z}^{i}$, then we can get all the synthetic images and collect them into $\mathcal{S}^{i}$.
@@ 3.1 Category Concepts Learning
The pre-trained datasets of diffusion models may have a distribution gap with the downstream classification benchmarks. Thus, it is hard to directly use category labels to guide the diffusion model to generate corresponding faithful images. Learning a more faithful concept for each category as the prompt for downstream generation is quite necessary. To achieve this, we followed the same learning strategy as (Wang et al., 2024). Specifically, there are two learnable parts: 1) Token embeddings: For the $i_{th}$ category, we set $n$ learnable concept tokens ($\{[V_{j}^{i}]\}_{j=1}^{n}$) 2) Low-rank matrixes: We insert some low-rank matrixes (Hu et al., 2021) into the pretrained diffusion U-Net. These matrixes are shared by all categories.
@@
where $x_{0}$ is the encoded latent of $X_{j}^{i}$, $\bar{\alpha_{t}}$ is a pre-defined parameter and $\epsilon$ is a Gaussian noise. The learning objective is:

### v2

# Inversion Circle Interpolation: Diffusion-based Image Augmentation for Data-scarce Classification

@@ Abstract
Data Augmentation (DA), i.e., synthesizing faithful and diverse samples to expand the original training set, is a prevalent and effective strategy to improve the performance of various data-scarce tasks. With the powerful image generation ability, diffusion-based DA has shown strong performance gains on different image classification benchmarks. In this paper, we analyze today’s diffusion-based DA methods, and argue that they cannot take account of both faithfulness and diversity, which are two critical keys for generating high-quality samples and boosting classification performance. To this end, we propose a novel Diffusion-based DA method: Diff-II. Specifically, it consists of three steps: 1) Category concepts learning: Learning concept embeddings for each category. 2) Inversion interpolation: Calculating the inversion for each image, and conducting circle interpolation for two randomly sampled inversions from the same category. 3) Two-stage denoising: Using different prompts to generate synthesized images in a coarse-to-fine manner. Extensive experiments on various data-scarce image classification tasks (e.g., few-shot, long-tailed, and out-of-distribution classification) have demonstrated its effectiveness over state-of-the-art diffusion-based DA methods.
Figure 1: Given training images, data augmentation aims to generate new faithful and diverse synthetic images. (a) These synthetic images are faithful but not diverse. (b) These synthetic images are diverse but not faithful. (c) These synthetic images are both faithful and diverse.
@@ 1 Introduction
Today’s visual recognition models can even outperform us humans with sufficient training samples. However, in many different real-world scenarios, it is not easy to collect adequate training data for some categories (i.e., data-scarce scenarios). For example, since the occurrence frequency of various categories in nature follows a long-tailed distribution, there are many rare categories with only limited samples [25, 34]. To mitigate this data scarcity issue, a prevalent and effective solution is Data Augmentation (DA). Based on an original training set with limited samples, DA aims to generate more synthetic samples to expand the training set.
@@
For DA methods, there are two critical indexes: faithfulness and diversity [29]. They can not only show the quality of synthesized samples, but also influence the final classification performance. More specifically, faithfulness indicates that the synthetic samples need to retain the characteristics of the corresponding category (cf., Figure 1 (a)), i.e., the faithfulness confirms that the model learns from correct category knowledge. Diversity indicates that the synthetic samples should have different contexts from the original training set and each other (cf., Figure 1 (b)), i.e., the diversity ensures that the model learns the invariable characteristics of the category by seeing diverse samples.
@@
Figure 2: a) Intra-category DA: Given a reference image (from the original set), it adds some noise and denoises with a prompt containing the same category concept (e.g., concept “[A]” for category A image). (b) Inter-category DA: Different from Intra-category DA, it denoises with a prompt containing a different category concept (e.g., concept “[B]” for category A image). (c) Ours: It first calculates the inversion for each image, and conducts random circle interpolation for two images of the same category. Then, it denoises in a two-stage manner with different prompts.
@@
With the photo-realistic image generation ability of today’s diffusion models [13, 23], a surge of diffusion-based DA methods has dominated the image classification task [^1] DA has been used in various image classification settings. In this paper, we focus on data-scarce scenarios (more discussions are in the Sec. 2).. Typically, diffusion-based DA methods reformulate image augmentation as an image editing task, which consists of two steps: 1) Noising Step: They first randomly sample an image from the original training set as a reference image and then add some noise to the reference image. 2) Denoising Step: They then gradually denoise this noisy reference image conditioned on a category-specific prompt. After the two steps, a new synthesized training image was generated. Following this framework, the pioneer diffusion-based DA work [12] directly uses a hand-crafted template containing the reference image’s category label as the prompt (i.e., intra-category denoising). These handcrafted prompts work well on general datasets with a broad spectrum of category concepts (e.g., CIFAR-10 [17]). However, these few words (with only category name) can not guide the diffusion models to generate images with specific and detailed characteristics, especially for datasets with fine-grained categories (e.g., Stanford Cars [16]).
@@
To further enhance the generalization ability, subsequent diffusion-based DA methods try to improve the quality of synthesized samples from the two key characteristics. Specifically, to improve faithfulness, [33] replace category labels with more fine-grained learned category concepts. As shown in Figure 2(a), they first learn a specific embedding “[A]” for “category A” bird, and then replace the fixed category name with the learned concept in the prompt. These learnable prompts can somewhat preserve fine-grained details for different categories. However, the fixed combination of a reference image and its corresponding category concept always results in similar synthetic samples (i.e., limited diversity). On the other side, to improve diversity, [36] use prompts containing different category concepts (e.g., “[B]”) from the reference image (i.e., inter-category denoising). This operation can generate images with “intermediate” semantics between two different categories. However, it inherently introduces another challenging problem to obtain an “accurate” soft label for each synthetic image, which affects faithfulness to some extent (cf., Figure 2(b)). Based on these above discussions, we can observe that: current state-of-the-art diffusion-based DA methods cannot take account of both faithfulness and diversity, which results in limited improvements on the generalization ability of downstream classifiers.
@@
In this paper, we propose a simple yet effective Diffusion-based Inversion Interpolation method: Diff-II, which can generate both faithful and diverse augmented images. As shown in Figure 2(c), Diff-II consists of three steps: 1) Category Concepts Learning: To generate faithful images, we learn a specific embedding for each category (e.g., “[A]” for category A) by reconstructing the images of the original training set. 2) Inversion Interpolation: To improve diversity while maintaining faithfulness, we calculate the inversion [^2] In the image generation field, the inversion refers to a latent representation that can be used to reconstruct the corresponding original image by the generative model. for each image of the original training set. Then, we sample two inversions from the same category and conduct interpolation. The interpolation result corresponds to a subsequent high-quality synthetic image. 3) Two-stage Denoising: To further improve the diversity, we prepare some suffixes [^3] More suffix examples are shown in the appendix. (e.g., “flying over water”, “standing on a tree branch”) that can summarize the high-frequency context patterns of the original training set. Then, we split the denoising process into two stages by timesteps. In the first stage, we denoise the interpolation results guided by a prompt containing the learned category concept and a randomly sampled suffix, e.g., “a photo of a [A] bird [suffix].” This design can inject perturbation into the early-timestep generation of context and finally contributes to diversity. In the second stage, we replace the prompt with “a photo of a [A] bird” to refine the character details of the category concept.
@@
Specifically, we first utilize some parameter-efficient fine-tuning methods (i.e., low-rank adaptation [14] and textual inversion [10]) to learn the concept embedding for each category. Then, we acquire the DDIM inversion [32] for each image from the original set conditioned on the learned concept. After that, we randomly sample two inversions within one category as one pair and conduct interpolation with random strengths. To align the distribution of interpolation results with standard normal distribution and get a larger interpolation space, we conduct random circle interpolation. Since each pair of images used for inversion interpolation belongs to the same concept, their interpolations will highly maintain the semantic consistency of this concept (i.e., it ensures faithfulness). Meanwhile, since both images have different contexts, the interpolations will produce an image with a new context (i.e., it ensures diversity). Finally, we set a split ratio to divide the whole denoising timesteps into two stages. In the first stage, we use a prompt containing the learned concept and a randomly sampled suffix3 to generate noisy images with diverse contexts (e.g., layout and gesture). In the second stage, we remove the suffix to refine the character details of the category concept. By adjusting the split ratio, we can control the trade-off between faithfulness and diversity. To extract all suffixes, we first utilize a pretrained vision-language model to extract all captions of the original training set, then leverage a large language model to summarize them into a few suffixes.
@@
  • We propose a unified view to analyze existing diffusion-based DA methods in scarce-data scenarios, we argue that they can not take account of both faithfulness and diversity well [^4] More quantitative and qualitative analyses are left in the appendix., which results in limited improvements.
  • We propose an effective diffusion-based DA method, that leverages the inversion circle interpolation and two-stage denoising to generate faithful and diverse images.
  • Comprehensive evaluation on three tasks has verified that our Diff-II can achieve effective data augmentation by generating high-quality samples.
@@ 2 Related Work
Diffusion-Based DA. With the emergence of diffusion models, diffusion-based DA [22, 15] becomes a popular solution. One DA setting [37, 2, 12, 42] is enhancing coarse-grained datasets. By tuning the diffusion model into the target domain [2] or leveraging the language model to generate general descriptions for characteristics [37]. The faithfulness of coarse-grained categories can be guaranteed.
Meanwhile, another more crucial setting is enhancing small-scale fine-grained datasets. Since it’s more difficult to generate fine-grained appearances, DA methods always need to extract detailed patterns from reference images (e.g., an extra concept learning step). For this setting, there are two main paradigms: 1) Latent perturbation [42, 9, 41] generate samples by perturbating latent codes in the latent space. Although these methods can generate diverse samples, due to the uncontrollable perturbation direction, the generated results sometimes deviate from the domain of the original dataset. Therefore, they heavily rely on extra over-sampling and filtering steps. 2) Image editing [12, 33, 8, 36] reformulate data augmentation as an image editing task [21]. However, due to the limitations of the editing paradigm, it’s difficult for them to take into account both the faithfulness and diversity of the synthetic samples. To tackle the above problem, our work proposes to generate new images by interpolating the inversions.
Interpolation-Based DA. For time series and text data, interpolation is a common approach for DA. Chen et al. [4] incorporate a two-stage interpolation in the hidden space to improve the text classification models. Oh et al. [24] propose to augment time-series data by interpolation on original data. In the computer vision community, there are some studies [6, 42] work on interpolation-based DA for image classification. However, how to combine the excellent generation ability of diffusion models and interpolation operation to obtain high-quality synthetic samples remains an important challenge. In this paper, we utilize inversion circle interpolation by considering the distribution requirement for the diffusion model.
@@ 3 Method
Figure 3: Pipeline of Diff-II. (1) Concept Learning: Learning accurate concepts for each category. (2) Inversion Interpolation: Calculating DDIM inversion for each image conditioned on the learned concept. Then, randomly sampling a pair and conducting random circle interpolation. (3) Two-stage Denoising: Denosing the interpolation results in a two-stage manner with different prompts.
@@
1) Category Concepts Learning: We first set $n$ learnable token embeddings for each category, and insert some learnable low-rank matrixes into the pretrained diffusion U-Net. By reconstructing the noised image of the original training set $\mathcal{O}$, we learn the accurate concept for each category. We denote the tokens of the $i_{th}$ category concept as $\{[V_{j}^{i}]\}_{j=1}^{n}$.
2) Inversion Interpolation: Take the $i_{th}$ category as an example, we form a prompt: “a photo of a $[V_{1}^{i}]$ $[V_{2}^{i}]$$[V_{n}^{i}]$ [metaclass]3”. The “[metaclass]” is the theme of the corresponding dataset, e.g. “bird” is the “[metaclass]” for dataset CUB [35]. Then, we calculate the DDIM inversion $I_{j}^{i}$ for each training sample $X_{j}^{i}\in\mathcal{O}^{i}$ conditioned on this prompt. All these inversions (from $\mathcal{O}^{i}$) made up the inversion pool $\mathcal{I}^{i}=\{I_{j}^{i}\}_{j=1}^{N_{i}}$ (cf., Sec. 3.2.1). After that, we randomly sample two inversions ($I_{a}^{i}$, $I_{b}^{i}$) from $\mathcal{I}^{i}$ and conduct random circle interpolation on this pair (cf., Sec. 3.2.2). The interpolation result is denoted as $Z$. We repeat the sampling and interpolation then collect all interpolation results into $\mathcal{Z}^{i}$.
3) Two-stage Denoising: Given an interpolation $Z\in\mathcal{Z}^{i}$, we denoise it as the initial noise in two stages. The main difference between the two stages is the prompt used. In the first stage, we use a suffixed prompt: “a photo of a $[V_{1}^{i}]$ $[V_{2}^{i}]$$[V_{n}^{i}]$ [metaclass] [suffix]”. In the second stage, we use a plain prompt: “a photo of a $[V_{1}^{i}]$ $[V_{2}^{i}]$$[V_{n}^{i}]$ [metaclass]”. Repeat two-stage denoising for each $Z\in\mathcal{Z}^{i}$, then we can get all the synthetic images and collect them into $\mathcal{S}^{i}$.
@@ 3.1 Category Concepts Learning
The pre-trained datasets of diffusion models may have a distribution gap with the downstream classification benchmarks. Thus, it is hard to directly use category labels to guide the diffusion model to generate corresponding faithful images. Learning a more faithful concept for each category as the prompt for downstream generation is quite necessary. To achieve this, we followed the same learning strategy as [36]. Specifically, there are two learnable parts: 1) Token embeddings: For the $i_{th}$ category, we set $n$ learnable concept tokens ($\{[V_{j}^{i}]\}_{j=1}^{n}$) 2) Low-rank matrixes: We insert some low-rank matrixes [14] into the pretrained diffusion U-Net. These matrixes are shared by all categories.
@@
where $x_{0}$ is the encoded latent of $X_{j}^{i}$, $\bar{\alpha}_{t}$ is a pre-defined parameter and $\epsilon$ is a Gaussian noise. Training objective is: